OpenStudy (anonymous):

In Gullivers travels by johnathan swift, gulliver went to the kingdom of Lilliput where he found that he was 12 time the height of the average lilliputian. 1.The lilliputians computed gullivers surface area to make him clothes. The amount of material needed for his suit is about how many times the amount of material needed for one of theirs. 2.The lillipyutains computed gullivers volume to determine how much food he would need. The amount gulliver required is about how many times the amount required by a lilliputian?

OpenStudy (anonymous):

How in the world do you find the surface area and volume of a human... Anyway, I assume, their width are the same. So, when his height is 12 times of the Liliputians, what would that make his surface area be? Assume it is a rectangle. I.E. Base x height

OpenStudy (anonymous):

Can you calculate how much larger is surface area would be now?

OpenStudy (anonymous):

 number of times surface area is larger = 24 / 2 = 12 When the length is increased by 12 times and the breath is kept constant, area would also be 12 times

OpenStudy (unklerhaukus):

each of Gullivers Height, Length , and Width scale by 12 A=H*L → 12H*12L V=H*L*W → 12H*12L*12W